OpenStudy (anonymous):

How do I find the integration of sin (pix^3)dx?

OpenStudy (dumbcow):

use numerical analysis. evaluate definite integral by estimating area under curve

OpenStudy (anonymous):

I know that, but how do I integrate sin with the pix^3? It's cos something isn't it? but I forgot how to integrate.

OpenStudy (dumbcow):

thats what i mean there is no way to integrate it using any methods (substitution, by parts...)

OpenStudy (anonymous):

ok, well, here is the real problem: it's a double integration: \[\int\limits_{0}^{9}\int\limits_{\sqrt{y}}^{3}\sin (pix^3) dxdy\]